A train starts from rest and travels 10 seconds with acceleration a=4-0.3t, then runs for 30 seconds at constant velocity and finally comes to rest by decelerating at 3 m/s^2. Determine the MAGNITUDE of the CONSTANT VELOCITY in m/s.
1
Expert's answer
2020-04-09T09:16:03-0400

"v_0=0" - A train starts from rest

"t=10s" - the time with acceleration

"a=4-0.3t"


"a_{avr}=\frac{a(t_0=0)+a(t_{10}=10)}{t_{10}-t_0}=\frac{0+1}{10}=0.1"


"v=v_0+\frac{a_{avr}t^2}{2}\\
v=0+\frac{0.1 \cdot 10^2}{2}=5" m/s


Answer: 5 m/s